When price of tea in a lacal cafe rises from Br.10to15 per cup, demand for coffee rises from 3000 cups to 5000 a day despite no change in coffe price
A. Determine cross price elasticity of demand
What kind of realtion exists between the two good?why?
1
Expert's answer
2021-08-18T07:03:45-0400

Solution:

A.). Cross price elasticity of demand = "=\frac{\%\; change\; in\; quantity\; of\; coffee\; demanded}{\%\; change\; in\; the \; price\; of\; tea}"


Cross price elasticity of demand = "=\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \div \frac{P_{2} -P_{1}}{(P_{2}+P_{1})/2 }"



= "\frac{5000 -3000}{(5000+3000)/2 } \div\frac{15 -10}{(15+10)/2 } = \frac{0.5}{0.4} = 1.25"


The cross-price elasticity of demand = 1.25


The two goods, that is coffee and tea are directly related and are substitutes since the cross-price elasticity of demand is greater than one.

They are two alternative products that are similar enough to be used in place of another. Therefore, when the price of one substitute, that is tea, goes up, the demand for the other substitute product, that is coffee, increases as consumers substitute one good, that is tea for another, that is coffee.